OpenStudy (anonymous):

Factor completely: xy + 4y − 2x − 8

OpenStudy (anonymous):

Factor completely: xy + 4y − 2x − 8 (x − 4)(y + 2) (x + 4)(y − 2) (x − 4)(x − 2) Prime

OpenStudy (ivancsc1996):

You do the same thing as in the question before.\[(x+4)(y-2)=xy-2x+4y-8=xy+4y-2x-8\]

OpenStudy (anonymous):

okay thank you!